Fermi National Laboratory


2006 Accelerator Update Archive

December 21, 2006 - January 3, 2007
- 13 stores provided luminosity to the experiments
- Store 5155 set a luminosity record with 248.2E30
- 3 quenches occurred with 2 ending stores
- NuMI turned off for absorber work

December 18, 2006-December 20, 2006
- Three stores provided 41 hours and 49 minutes of luminosity
- I- Source trips continue
- NuMI BPM and horn problems
- Linac Debuncher water skid repaired

December 15, 2006-December 18, 2006
- Three store provided 55 hours and 42 minutes of luminosity
- Electron cooling off due to LCW leak
- MiniBooNE resumes running with horn
- Meson experiment starts and ends its run this weekend
- Store 5128 sets second highest luminosity
- Store 5128 aborted due to lead trip

December 13, 2006-December 15, 2006
- Three stores provided 14 hours and 51 minutes of luminosity
- MI abort kicker power supply repaired
- MiniBooNE LCW pump repaired
- Store 5123 lost due to TeV quench
- TeV quenched during dry squeeze preparation
- Pbar magnet's hose breaks
- NuMI has kicker problem
- MCenter T959 experiment coming online today

December 11, 2006-December 13, 2006
- Three stores provided 39 hours and 30 minutes of luminosity
- Pbar's target stopped rotating
- Four MI-10 vacuum pumps trip off
- MI kicker problems
- MiniBooNE off due to LCW pump problem

December 8, 2006-December 11, 2006
- Three stores provided 59 hours and 12 minutes of luminosity
- TeV and MI suffer from VCB problems
- Recycler has kicker trouble
- MI has kicker trouble
- MI had MI beam fall out during the weekend - investigations continue
- Pbar has LCW leak

December 6, 2006-December 8, 2006
- Three stores provided 26 hours and 13 minutes of luminosity
- MI suffers from NuMI beam fall out
- Switchyard 120 gets OK to take beam
- Store 5008 lost due to power supply trip
- I- Source trips; H- brought online
- MiniBooNE begins horn off running mode

December 4, 2006-December 6, 2006
- Three stores provided 38 hours and 53 minutes of luminosity
- MI vacuum problems
- Booster repeater and BOM trouble
- NuMI suffers Intermittent beam loss when MI stashes
- Linac Debuncher water skid trip

December 1, 2006-December 4, 2006
- Four stores provided 45 hours and 40 minutes of luminosity
- NuMI pulsing spare horn
- Two stores lost
- Accumulator LCW leak causes ground fault
- NuMI off due to BPM problem
- MI low level RF problems

November 29, 2006-December 1, 2006
- Two stores provided 40 hours and 45 minutes of luminosity
- ComEd power glitch causes equipment trips, store continued
- Booster and Linac repair RF stations
- NuMI and MiniBooNE took beam

November 24, 2006-November 29, 2006
- Four stores provided 95 hours and 42 minutes of luminosity
- CHL had cold box problems
- Recycler had instrumentation problems
- Un-established store lost due to TeV dipole power supply failure
- TeV A1 cryo system had vacuum trouble
- MI-20 fire alarm caused by lighting ballast failure - stash lost

November 17, 2006-November 20, 2006
- Three stores provided 39 hours and 26 minutes of luminosity
- Preaccelerator switched to I- Source
- Recycler has cooling problems
- Pbar target air-pressure pump shaft breaks
- Store 5082 hits third largest luminosity with 210.12E30

November 15, 2006-November 17, 2006
- Two stores provided 6 hours and 47 minutes of luminosity
- Recycler had stashing problems
- TeV quench due to orbit drift
- Pbar had beam position problems
- TeV BPM trouble in F sector
- Power glitch: store 5074 lost, Linac and Booster go down
- Machines have trouble recovering from feeder 47 glitch

November 13, 2006-November 15, 2006
- Two stores provided 45 hours and 53 minutes of luminosity
- Pbar LCW leak found in Transport
- Booster kicker problems
- NuMI off for chiller work
- PBar lithium lens trouble
- Recycler damper work
- TeV records the Kuril Island earthquake

November 10, 2006-November 13, 2006
- Three stores provided 55 hours and 34 minutes of luminosity
- Recycler has emittance problem
- Cable for Booster chopper fails
- Pbar LCW leak

November 8, 2006-November 10, 2006
- Two store provided 27 hours and 58 minutes of luminosity
- Recycler looses 150E10 of stash
- Lithium lens trips off
- Separator spark causes store 5055 to quench
- Quadrupole trip causes store 5057 to quench
- Linac Klystron Debuncher and RF problems

November 6, 2006-November 8, 2006
- One store provided 20 hours and 26 minutes of luminosity
- 400 MeV Debuncher cavity temperature problems
- Recyclers looses 100.0E10 antiprotons

November 3, 2006-November 6, 2006
- TeV E4 vacuum certified
- MI RF cavity LCW leak
- I- Source trips
- Pbar resumes stacking
- NuMI and MiniBooNE resumes taking beam
- Recycler resumes stashing
- NuMI kicker failure
- Linac Klystron Debuncher problems

November 1, 2006-November 3, 2006
- TeV vacuum work continues
- Decomposed O ring found at E4

October 27, 2006-October 30, 2006
- Three stores provided 50 hours and 50 minutes of luminosity
- NuMI held off
- TeV quench - store lost before established
- TeV quench - store 5037 lost

October 25, 2006-October 27, 2006
- ComEd glitch trips of Booster
- Preaccelerator switches to I- Source
- Store 5028 aborted due to safety system trip
- A PLC problem halts stashing

October 23, 2006-October 25, 2006
- Three stores provided 14 hours and 50 minutes of luminosity
- Bad cable at TeV B4 cryo system causes two quenches
- Safety system glitch causes store abort
- D17 separator spark

October 20, 2006-October 23, 2006
- Booster Kicker problems
- H- Source current drop
- Accumulator ground fault
- MI RF vacuum burst
- Tevatron beam established

October 18, 2006-October 20, 2006
- TeV and Pbar magnet replacement continues
- Pbar verified stacking
- TeV sector B3 cold compressor tripped off
- Pbar magnet might have ground fault
- TeV cooldown begins on the evening of 10/21/06

October 13, 2006-October 16, 2006
- The Tevatron magnet replacement continues
- NuMI resumed taking beam, but problems continued
- Booster power supply troubles
- Carbon debris found in TeV beam pipe
- LCW leak found in Pbar's AP2 transport line
- Pbar runs beam to dump

October 11, 2006-October 13, 2006
- The Tevatron shutdown to replace a D3 magnet continues
- The I- Source has a cesium problem
- NuMI magnet LCW leak repaired
- NuMI power supply (HV101A) fails
- Safety System testing.

October 9, 2006-October 11, 2006
- One store provided 18 hours and 19 minutes of luminosity
- Vacuum burst causes Recycler to loose part of stash
- Power glitch quenches TeV, looses the store, ground faults magnet, trips Linac, Booster, and many systems - all caused by a mouse
- H- Source shorts out (not caused by a mouse)
- TeV off for a week

October 6, 2006-October 9, 2006
- Three stores provided 63 hours and 28 minutes of luminosity
- 10/6/06, Pbar sets New Stacking Record of 20.56 mA/hr
- P1 power supply causing NuMI trips
- Accumulator RF cavity fills with water
- H- Source problems
- Booster kicker needs reterminating

October 2, 2006-October 6, 2006
- Two stores provided 22 hours and 56 minutes of luminosity
- Pbar set a stacking record on Friday with a rate of 20.53 mA per hour
- Store 5003 lost after only 28 minutes

September 29, 2006-October 2, 2006
- Four stores provided 32 hours and 34 minutes of luminosity
- MI power supply ground fault
- Store 4990 quenched due to possible separator spark
- MI power supply causes NuMI trips

September 27, 2006-September 29, 2006
- One store provided 4 hours and 10 minutes of luminosity
- TeV suffers three quenches
- TeV kicker replaced
- Pbar Lens work completed

September 25, 2006-September 27, 2006
- One store provided 11 hours and 31 minutes of luminosity
- TeV suffers from quenches and aborts
- Antiproton target repaired
- TeV kicker magnet replacement begin 9/28/06

September 22, 2006-September 25, 2006
- Three stores provided 67 hours and 57 minutes of luminosity
- Store 4967 aborted due to controls link problem
- Pbar target stops rotating
- Linac RF station holds off beam
- TEL2 studies

September 13, 2006-September 15, 2006
- Two stores provided 18 hours and 9 minutes of luminosity
- Store 4952 quenched

September 11, 2006-September 13, 2006
- Two stores provided 45 hours and 36 minutes of luminosity
- Recycler looses antiprotons due power supply trip
- NuMI resumes taking beam
- TeV sector E48 separator spark

September 8, 2006-September 11, 2006
- Three stores provided 54 hours and 2 minutes of luminosity
- Store 4942 lost due to quench
- TeV abort during dry squeeze
- Stores 4947 and 4949 set luminosity records
- Booster kicker problems
- Serious controls system problems resolved by experts
- NuMI starting up today

August 9, 2006-August 11, 2006
- Two stores provided 42 hours and 30 minutes of luminosity.

August 4, 2006-August 9, 2006
- Three stores provided 32 hours and 30 minutes of luminosity.
- Water leaks and bad pump motors.

August 2, 2006-August 4, 2006
- Fermi lower power usage per ComEd request
- Half of stash lost due to lightning strike
- Store 4868 lost due to lightning strike
- Booster quad trip hold off beam
- Stash lost due to Power supply problem
- TeV shot setup aborted
- Machine Reports

July 31, 2006-August 2, 2006
- Two stores provided 31 hours and 54 minutes of luminosity
- H- source failing
- Pbar problems continue
- High temperatures causing problems

July 28, 2006-July 31, 2006
- Two stores provided 54 hours and 4 minutes of luminosity
- NuMI Lambertsons trip
- Divers continue to clean strainers
- Pbar problems: Overthruster, BPMs, ground faults, failed magnet
- High temperatures affect every machine
- TeV D48 polarity problems

July 26, 2006-July 28, 2006
- Two stores provided 42 hours and 51 minutes of luminosity
- NuMI resumes taking beam
- NuMI suffers many trips
- Store 4859: second best luminosity with 171.2E30
- Booster RF problems
- Machine Reports

July 24, 2006-July 26, 2006
- Two stores provided 25 hours and 19 minutes of luminosity
- Controls system problems
- Feeder 49 ground fault causes power glitch
- TeV quench due to kicker failure
- NuMI to start taking beam on Wednesday afternoon (7-26-06)

July 21, 2006-July 24, 2006
- Three stores provided 63 hours and 8 minutes of luminosity
- H- Source problems
- MTest beam disabled
- Booster WAPS trips
- ARF1 trouble
- Recycler has electron cooling trouble

July 19, 2006-July 21, 2006
- Three stores provided 26 hours and 33 minutes of luminosity
- Recycler low-level RF problem
- Booster heat exchanger clogs
- Store 4839 aborted due to RF trip
- Store 4841 aborted due to ramp trip
- Machine Reports

July 17, 2006-July 19, 2006
- Booster has BRF17 trouble
- Linac's line voltage sags
- Stash lost
- MI has BPM problems
- LCW temperatures rise

July 14, 2006-July 17, 2006
- Four stores provided 47 hours and 36 minutes of luminosity
- Booster has RF problems
- FESS and LCW technicians clean pond strainers
- Multiwire stuck in beam line
- Store 4829 aborted
- Pbar PLC fails
- Store 4835 quenched
- Linac Q5BULK supply problem

July 12, 2006-July 14, 2006
- Three stores provided 25 hours and 3 minutes of luminosity
- NuMI repairs progressing
- Booster quadrupole problems
- Two stores aborted due to B4 vacuum trouble
- TeV A1 heat exchanger plugged
- Machine Reports

July 10, 2006-July 12, 2006
- Two stores provided 33 hours and 4 minutes of luminosity
- NuMI still off
- Accumulator has quadrupole problem
- Store 4821 aborted
- D0 might require 2 shift access Thursday

July 7, 2006-July 10, 2006
- Three stores provided 51 hours and 58 minutes of luminosity
- NuMI Horn still off
- TeV RF station trips off
- Booster kicker and chopper trips
- Store 4816 terminated to save lose due to potential quench
- Bad PLC causes Pbar to unstack

July 6, 2006-July 7, 2006
- One store provided 24 hours of luminosity
- A0 pond strainer cleaned
- NuMI off
- Machine Reports

July 5, 2006-July 6, 2006
- One store provided 24 hours of luminosity
- NuMI off due to horn cooling problem
- Booster access to repair five PAs
- A0 cold compressors trip

June 30, 2006-July 5, 2006
- Five stores provided ~ 86 minutes and 43 minutes of luminosity
- NuMI off due to horn cooling problems
- Store 4805 aborted
- TeV quench in between stores
- Recycler lost small portion of stash
- MI-40 power supply shorted

June 29, 2006-June 30, 2006
- Lithium Lens trips off
- D4 wet engine repair
- NuMI trips
- Proton abort kicker trips
- KRF4 trips off
- Machine Reports

June 28, 2006-June 29, 2006
- Two stores provided 19 hours and 56 minutes of luminosity
- LLRF problems halt stash
- Power supply trips hinder Switchyard
- NuMI horn trips
- CDF's solenoid tripped off

June 27, 2006-June 28, 2006
- Two stores provided 19 hours and 4 minutes of luminosity
- NuMI trips
- Water in lower Linac
- KRF2 problems
- A2 LCW leak
- MiniBooNE Horn trips

June 26, 2006-June 27, 2006
- Two store provided 21 hours and 8 minutes of luminosity
- Linac has Klystron problems
- Attempt to stash failed
- Nitrogen Dewar develops leak

June 23, 2006-June 26, 2006
- Three stores provided 58 hours and 55 minutes of luminosity
- A Pbar tuning program causes problems
- TeV suffered quench at sectors F2 and F3
- MTest begins taking beam
- Booster kicker problem
- H- Source tripped off

June 22, 2006-June 23, 2006
- One store provided 4 hours and 20 minutes of luminosity
- TeV orbit stabilization program problem understood
- Switchyard beam to F17
- Machine reports

June 21, 2006-June 22, 2006
- One store provided two minutes of luminosity
- Booster chopper problems
- D0 access completed
- Store 4787 quenched

June 20, 2006-June 21, 2006
- Two stores provided 12 hours and 39 minutes of luminosity
- Store 4782: initial average luminosity of 111.37E30
- Linac quadrupole magnet shorted
- MiniBooNE Horn power supply repaired
- Switchyard problems
- Store 4782 aborted
- D0 begins its 16-hour access

June 19, 2006-June 20, 2006
- One store provided 24 hours of luminosity
- Booster converts kicker to notcher
- Pelletron temperature problems continue
- D0 16-hour access begins tomorrow morning

June 16, 2006-June 19, 2006
- Four stores provided 43 hours and 8 minutes of luminosity
- Store 4781: initial average luminosity of 103.9E30
- Store 4774 aborted
- 400 MeV Debuncher water skid trips reappear
- MiniBooNE Horn power supply trip
- Power supply LCW leak holds off Pbar

June 15, 2006-June 16, 2006
- One store provided 18 hours and 53 minutes of luminosity
- Linac RF station holds off all beam
- Vacuum problem halts NuMI beam
- Machine Reports

June 14, 2006-June 15, 2006
- One store provided 5 hours and 59 minutes of luminosity
- NuMI off due to Lambertson problem
- Store 4760 accidentally aborted
- TeV quenches
- Linac RF station halts beam

June 13, 2006-June 14, 2006
- Two stores provided 20 hours and 35 minutes of luminosity
- TeV separator's wrong polarity distorts store luminosity
- D0 powers down solenoid for calibration
- CDF's silicon detector off

June 12, 2006-June 13, 2006
- One store provided 8 hours and 14 minutes of luminosity
- TeV quenched
- NuMI Horn ground fault resolved
- Multiple Klystron Debuncher cavity water skid trips

June 9, 2006-June 12, 2006
- TeV quenches twice
- Klystron Debuncher water skid problems continue
- TeV store 4745 established - First store since startup
- Pbar vacuum trouble
- NuMI target cooling problems
- Recycler stash lost
- MI power supply ground fault trips

June 8, 2006-June 9, 2006
- Pbar's ground fault resolved
- Booster still suffering from Debuncher water skid trips
- TeV quenched at F1
- Machine Reports

June 7, 2006-June 8, 2006
- Recycler stashed
- A:QT interferes with stacking
- Booster suffers from kicker and Debuncher cavity water problems
- Machine Reports

June 6, 2006-June 7, 2006
- TeV loads 12x0 study store
- Pbar stacks antiprotons
- D0 begins two shift access
- Recycler stashed antiprotons
- Klystron Debuncher trips
- Machine Reports

June 5, 2006-June 6, 2006
- MI's orbit shift discovered
- Pbar begins stacking
- MiniBooNE beam established
- MI conducts mixed mode stacking study
- Machine Reports

June 2, 2006-June 5, 2006
- NuMI ran well despite numerous trips
- Pbar ran reverse protons
- TeV quenches twice
- MI suffers from orbit shift
- D0 to make 16-hour access on Tuesday or Wednesday
- Machine Reports

June 1, 2006-June 2, 2006
- Startup
- NuMI issued beam permit
- Beam to NuMI target
- Beam in TeV
- Beam to Pbar target
- Switchyard safety system test completed
- Machine Reports

May 31, 2006-June 1, 2006
- Startup
- TeV receives permission to run beam
- Machine Reports

May 30, 2006-May 31, 2006
- Startup
- F4 wet engine broke connecting rod
- Cross Gallery attacked
- D0 powered solenoid and magnets
- Machine Reports

May 26, 2006-May 30, 2006
- TeV established 980 GeV ramp
- TeV suffers first quench
- Lightning and wind causes problems
- Linac gallery attacked
- Machine reports

May 25, 2006-May 26, 2006
- Startup
- Bad magnet found in P3 line
- Machine Reports

May 24, 2006-May 25, 2006
- Startup
- CDF powered solenoids
- Operations powered the TeV bulk power supplies
- Machine Reports

May 23, 2006-May 24, 2006
- Startup
- Beam established to MiniBooNE
- MI and Recycler succeeded in circulating beam
- Machine Reports

May 22, 2006-May 23, 2006
- Startup
- 8 GeV, MI, Recycler, Pelletron, MiniBooNE given permission to run
- Beam established to Booster and MI
- Machine Reports

May 19, 2006-May 22, 2006
- Shutdown ends and startup begins
- Linac and Booster have beam
- NTF studies
- Ground faults
- Machine Reports

May 12, 2006-May 19, 2006
- The shutdown has come to an end and the startup has begun.

May 5, 2006-May 12, 2006
- Shutdown Summary

April 28, 2006-May 5, 2006
- Shutdown Summary

April 21, 2006-April 28, 2006
- Shutdown Summary

April 14, 2006-April 21, 2006
- Shutdown Summary

April 7, 2006-April 14, 2006
- Shutdown Summary

March 31, 2006-April 7, 2006
- Shutdown work activities - the 6th week.

March 24, 2006-March 31, 2006
- Shutdown Summary

March 17, 2006-March 24, 2006
- Shutdown Summary

March 10, 2006-March 17, 2006
- Linac resumed operation on 3-9-06

March 3, 2006-March 10, 2006
- Linac resumed operation on 3-9-06

February 24, 2006-March 3, 2006
- Machine installations, upgrades, and repairs begin.

February 24, 2006-February 27, 2006
- Operations turned off all the machines at 0001 hours on Sunday, February 26.
- The shutdown offically began at 0000 hours on Monday, February 27.
- The Fermi Accelerator Complex has begun its twelve week shutdown for  new equipment installation, upgrades, and repairs.  An Accelerator  Update will come out every Friday covering the progress of the shutdown.

February 22, 2006-February 24, 2006
- Two stores provided 3 hours and 15 minutes of luminosity.
- TeV store 4666 quenched.
- TeV cryo system warm up begins early.
- Recycler and Main Injector conduct boomerang tests.
- Twelve-week shutdown begins 2/27/06.

February 20, 2006-February 22, 2006
- Two stores provided 22 hours and 51 minutes of luminosity.
- TeV quenched with no store.
- Store 4658 lost due to vacuum trouble.

February 17, 2006-February 20, 2006
- Three stores provided 47 hours and 38 minutes of luminosity.
- TeV quench during squeeze.
- One antiproton transfer lost for store 4653.

February 15, 2006-February 17, 2006
- One special 1x1 store provided 40 hours and 20 minutes of low luminosity.
- Booster suffered from a vacuum leak and RF problems.
- Linac also had RF trouble.

February 13, 2006-February 15, 2006
- Two stores provided 41 hours and 26 minutes of luminosity
- MCenter running without analyzing magnet
- Store 4642 quenched

February 10, 2006-February 13, 2006
- Three stores provided 59 hours and 59 minutes of luminosity.
- TeV quenches twice
- Preaccelerator switched to H-
- Recycler losses part of stash

February 8, 2006-February 10, 2006
- One store plus an existing store provided 41 hours and 28 minutes of luminosity
- NuMI beam problems continued
- Pbar and MCenter have LCW leaks

February 6, 2006-February 8, 2006
- Two stores provided 39 hours and 14 minutes of luminosity
- The Antiproton Source made a stacking record with 17.4 mA/hr
- Pelletron problem resolved

February 3, 2006-February 6, 2006
- Three stores plus an existing store provided 52 hours and 12 minutes of luminosity
- TEL trips resolved
- Store 4624 aborted due to a quench
- Accumulator quadrupole ground fault

February 1, 2006-February 3, 2006
- Two stores provided 34 hours and 5 minutes of luminosity
- Pbar Lens repaired
- F17B3 power supply problem holds off beam

January 30, 2006-February 1, 2006
- One low luminosity store provided 30 hours and 9 minutes of luminosity
- TeV sector C3 Wet Engine work
- Tevatron Electron Lens trips
- Pbar Lithium Lens problems

January 27, 2006-January 30, 2006
- During the 72 hour period operations established two stores that provided 50 hours and 30 minutes of luminosity.

January 25, 2006-January 27, 2006
- Working towards a Tevatron store.

January 23, 2006-January 25, 2006
- Tevatron startup in progress.

January 20, 2006-January 23, 2006
- Tevatron cooldown is underway.

January 18, 2006-January 20, 2006
- Tevatron still down for vacuum repairs.

January 13, 2006-January 18, 2006
- During the 120 hour period operations established 0 stores, that along with a previous store provided the experiments with 11 hours and 45 minutes of luminosity.
- Tevatron quench causes vaccum problems, that will could keep the Tev off for 2 weeks.

January 11, 2006-January 13, 2006
- During the 48 hour period operations established two stores that along with a previous store provided the experiments with 35 hours and 15 minutes of luminosity.

January 9, 2006-January 11, 2006
- During the 48 hour period operations established one stores that along with a previous store provided the experiments with 42 hours of luminosity.

January 6, 2006-January 9, 2006
- During the 72 hour period operations established two stores that along with a previous store provided the experiments with 61 hours and 15 minutes of luminosity.
- Booster RF problems caused lower intensities at times.

January 4, 2006-January 6, 2006
- During the 48 hour period operations established one store that provided the experiments with 4 hours and 30 minutes of luminosity.
- Kicker and Vacuum problems make for a bad 2 days.
- New recycler stash record of 370.1 E30 set on Thursday evening.
- New Tev Initial luminosity record of 171.85 E30 set on friday owl shift.

January 1, 2006-January 4, 2006
- During the 80 hour period operations established one store that along with a previous store provided the experiments with 58 hours and 45 minutes of luminosity.
- New recycler stash record of 350.6 E30 set on Tuesday afternoon.
- Kicker and Collimator problems cause unwanted downtime.

[an error occurred while processing this directive]

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.


Return to Current Accelerator Update



last modified 12/28/2011    email Fermilab